OpenStudy (chaotic_butterflies):

sketch the graph of each equation: x^2 + y^2 = 36

OpenStudy (chaotic_butterflies):

How should I go about graphing this? Should I get y on one side and create a chart of points?

OpenStudy (chaotic_butterflies):

@inkyvoyd could you help me?

OpenStudy (pokieokie):

What grade are you in?

OpenStudy (chaotic_butterflies):

11th

OpenStudy (pokieokie):

Oh Okay XD I would help, but your way up there!! I am still trailin 9th going into 10th

OpenStudy (pokieokie):

I apologize, forgive me

OpenStudy (chaotic_butterflies):

I'm taking a 12th/college level course

OpenStudy (chaotic_butterflies):

You're fine. @PokieOkie

OpenStudy (pokieokie):

Alright somebody should help you! Take care of yourself! Have a wonderfully blessed and safe week!

OpenStudy (chaotic_butterflies):

@sammixboo could you help?

satellite73 (satellite73):

\[x^2 + y^2 = 36\]

satellite73 (satellite73):

standard form for a circle with center \((h,k)\) and radius \(r\) is \[(x-h)^2+(y-k)^2=r^2\]by pythagoas

satellite73 (satellite73):

in your case \((h,k)=(0,0)\) so the center is at the origin

OpenStudy (chaotic_butterflies):

Oh.. I didn't even think about it like that

satellite73 (satellite73):

and here \(r^2=36\)

OpenStudy (chaotic_butterflies):

So what would x and y be?

OpenStudy (chaotic_butterflies):

just variables?

OpenStudy (chaotic_butterflies):

or a given point?

jimthompson5910 (jim_thompson5910):

One way is to isolate y to get two separate equations. Graph each equation using a T chart. After you've got your graphs set up, you can use something like desmos to check your work https://www.desmos.com/calculator I recommend doing the work by hand first, then use technology to check afterward

jimthompson5910 (jim_thompson5910):

or you can use @satellite73 's method to recognize that it's a form of a circle. And use the properties of a circle (center and radius) to plot the graph

OpenStudy (chaotic_butterflies):

That's how I thought about approaching it, but @satellite73 's approach is what I believe my book is wanting me to do

jimthompson5910 (jim_thompson5910):

ok then yeah use the second method